At a certain company the ratio of male to female employees is 3 to 4.If there are 5 more female employees than male employees, how many male employees does the company have
Answer by stanbon(75887)

At a certain company the ratio of male to female employees is 3 to 4.If there are 5 more female employees than male employees, how many male employees does the company have
=========
# of male = 3x
# of female = 4x
Equation:
4x-3x = 5
x = 5
---
Ans: # of male employees = 3x = 3*5 = 15
-----------
Cheers,
Stan H.
-------------